# QCRD: Quality-guided Contrastive Rationale Distillation for Large Language Models

### v1

@@ Abstract
Deploying large language models (LLMs) poses challenges in terms of resource limitations and inference efficiency. To address these challenges, recent research has focused on using smaller task-specific language models, which are enhanced by distilling the knowledge rationales generated by LLMs. However, previous works mostly emphasize the effectiveness of positive knowledge, while overlooking the knowledge noise and the exploration of negative knowledge. In this paper, we first propose a general approach called quality-guided contrastive rationale distillation for reasoning capacity learning, considering contrastive learning perspectives. For the learning of positive knowledge, we collect positive rationales through self-consistency to denoise the LLM rationales generated by temperature sampling. For the negative knowledge distillation, we generate negative rationales using temperature sampling for the iteration-before smaller language models themselves. Finally, a contrastive loss is designed to better distill the positive and negative rationales into the smaller language model, where an online-update discriminator is used to judge the qualities of rationales and assign weights for better optimizing the training process. Through extensive experiments on multiple reasoning tasks, we demonstrate that our method consistently outperforms the previous distillation methods and produces higher-quality rationales.
@@ 1 Introduction
The reasoning capabilities of large language models (LLMs) have been observed to scale with their model sizes, while necessitating substantial memory and computing resources Chowdhery et al. (2023); Wei et al. (2022a). As such, efficient model compression is crucial in the deployment of LLMs, especially on resource-limited devices or platforms. In recent years, knowledge distillation from an LLM to a smaller language model has emerged as a powerful and promising technique for model compression Hinton et al. (2015); Phuong and Lampert (2019). However, it is still open how to best close the performance gap with respect to LLMs on complex reasoning tasks Zelikman et al. (2022).
In this regard, it has more recently been shown that adding explanation-augmented prompts, especially, Chain-of-Thought (CoT) Wei et al. (2022b), can enable LLMs to generate reasonable explanations to justify the reasoning outcomes Li et al. (2022). In addition, distilling these rationales into smaller language models has been demonstrated to effectively improve the overall performance Hsieh et al. (2023); Li et al. (2022). Specifically, distilling Step-by-Step (DSS) Hsieh et al. (2023) was proposed as an innovative CoT distillation approach, which employed rationales from an LLM to guide a smaller language model under a multi-task learning setting. The approach involved training the smaller language model simultaneously on both label prediction and rationale generation tasks, effectively leveraging their mutual benefits.
@@
The essence of distilling rationales is to guide the model in learning additional knowledge related to the labels, and the knowledge can be concluded into two class: positive and negative. Previous works, although effective, still suffer from some drawbacks. On the one hand, the positive knowledge for distillation is noisy. Methods Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022) treat the rationales generated by LLMs as golden answers and aim to minimize the gap between these rationales and those generated by smaller language models. However, despite LLM’s powerful zero-shot/few-shot abilities, it may occasionally produce incorrect reasoning steps, leading to erroneous rationales/answers. Such erroneous rationales may degrade the reasoning performance of the distilled smaller language models. On the other hand, generating negative rationales and incorporating them into CoT distillation remain understudied, while negative knowledge has early proved to be constructive and effective for models.
@@
To this end, we propose a general method, named Quality guided Contrastive Rationale Distillation (QCRD), to guide the knowledge distillation to a smaller language model from a contrastive learning perspective. The comparison between the previous methods and our proposed QCRD is illustrated in Fig. 1. Specifically, the previous methods focus on the alignment between the rationale of the student model and the corresponding one of the teacher model, while our proposed QCRD aligns the student’s distribution and contrastive knowledge distribution with different sampled rationales. The core design of QCRD is to generate a diverse set of contrastive rationales and efficiently distill them into the student model. For the positive part, to ensure the qualities of positive rationales, we first prompt the LLM and sample the output to generate multi-round rationales for each input question. We then apply the self-consistence to denoise the rationale set and split it into positive samples and negative samples. For the negative part, we employ a self-play strategy Silver et al. (2018) during the training to generate low-quality rationales from previous iterations of the smaller language model and treat them as negative samples. Finally, for better knowledge learning, we present a contrastive loss to distill both the positive and negative rationales into the smaller language model. And a discriminator is adopted to assess the qualities of the rationales and assign them appropriate weights to optimize the training process across the datasets. To demonstrate the superiority of QCRD, we conduct comprehensive experiments with two smaller types of T5 models Raffel et al. (2020), i.e., T5-base (220M parameters) and T5-small (60M parameters), on four popular datasets, followed by detailed analysis and discussion.
Our main contributions of this paper can be summarized below.
@@
Figure 2: Illustration of the proposed quality guided contrastive rationale distillation (QCRD) for distilling contrastive knowledge from teacher models into the student model, where CE denotes the cross entropy.
@@ 2 Related Work
Our work utilizes the emerging reasoning capabilities of current LLMs to distill task-specific knowledge into smaller specialist models. We present a brief overview of previous studies across four areas related to our work in the following.
Knowledge distillation from LLMs. Knowledge distillation (KD) is a highly effective technique for transferring knowledge from larger, more capable teacher models to smaller student models that are more suitable for practical applications Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022). The technique can be generally classified into two different categories, namely, Black-box KD and White-box KD. White-box KD provides access to the teacher’s parameters, enabling the student LM to gain a deeper understanding of the teacher LLM’s internal structure and knowledge representations, while only the teacher’s predictions are accessible in Black-box KD. Both of Black-box and White-box KDs have shown promising results in fine-tuning small models on the prompt response pairs generated by LLMs Zhu et al. (2023). In this paper, we hypothesize that only the prediction labels and rationales generated by LLMs are accessible.
Multi-task learning with LLM generated rationales. Current LLMs have already shown their capabilities to generate high-quality reasoning steps, resulting in explanations of their predictions Kojima et al. (2022), and these explanations have been found to be valuable additional knowledge for fine-tuning smaller models Hsieh et al. (2023). In order to empower smaller models with robust reasoning abilities and the capacity to generate explanations, a multi-task learning framework is commonly employed. This framework enables the integration of strong reasoning capabilities into smaller models while maintaining their original functionality. However, previous approaches have applied a single loss form, which only focuses on aligning the output of the small model with that of the LLM Hsieh et al. (2023); Magister et al. (2022).
Self-consistency of LLMs. The self-consistency of LLMs refers to the capacity to maintain coherent and rational reasoning during input processing. Based on the intuition that complex reasoning tasks typically admit multiple reasoning paths that reach a correct answer, the self-consistency can improve the LLMs’ reasoning performance by integratedly sampling CoT outputs and choosing the most consistent answer Stanovich and West (1991). The self-consistency has proven to be an effective technique for denoising datasets and selecting positive and negative samples Wang et al. (2022).
Contrastive learning for LLMs. Contrastive learning has demonstrated its efficiency across diverse domains, encompassing computer vision, natural language processing, and recommendation systems Jaiswal et al. (2020); Le-Khac et al. (2020). Notably, the application of contrastive learning to LLMs has emerged as a recent development, highlighting the effectiveness of incorporating negative samples and showing promising outcomes Li et al. (2024). However, to the best of our knowledge, the application of contrastive learning in CoT distillation has not been explored thus far.
@@ 3 Methodology
In this paper, we first propose a general contrastive CoT distillation approach, i.e., quality guided contrastive rationale distillation (QCRD), for training smaller models by distilling contrastive knowledge from teacher models. Our approach leverages the reasoning capabilities and self-consistency of LLMs, as well as the self-supervised ability of smaller models themselves. As illustrated in Fig. 2, our approach consists of the following three parts. (1) Following the method described in Hsieh et al. (2023), we apply a multi-task framework for the supervised training of the student model, where the label task aligns the prediction label with the ground truth and the rationale task learns the knowledge from the generated rationales, as can be seen in Fig. 2a. (2) As displayed in Fig. 2b, we design a general approach to generate contrastive knowledge. Specifically, for the positive knowledge, we prompt the LLM to generate multi-round output labels and corresponding rationales using temperature sampling. Collected rationale samples are denoised based on the ground truth or the majority of the self-consistency results. For the negative knowledge, we consider the remaining noised rationales from the LLM as negative samples. Moreover, we generate additional lower-quality rationales by iterating the smaller model itself. These lower-quality rationales are thought of as negative samples. (3) As shown in in Fig. 2c, for better knowledge learning from rationales, We design a quality-guided contrastive learning strategy, where a contrastive loss is applied with the guidance of an online-updated discriminator to distinguish between positive and negative samples and assign them quality scores.
3.1 Multi-task framework for the student model
Previous works have already demonstrated the advantages of the multi-task framework. We follow the similar framework in Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022), where the label prediction task and rationale task are applied to the training of the smaller language model. Specifically, we use different prompts to enforce the smaller language model to generate different types of output. Given an input question, for the label prediction task, the smaller language model outputs the prediction label while for the rationale task, it outputs the explanation about the prediction label.
@@ 3.2 Generation of contrastive knowledge
In what follows, we introduce the generation of the contrastive knowledge for dsitillation in detailed.
3.2.1 Extract diverse rationales from LLMs
We use Chain-of-Thought (CoT) prompting Wei et al. (2022b) to elicit and extract rationales from LLMs. As illustrated in Fig. 3, the LLM is provided with few examples to follow the output format. Instead of only generating one output for each input, we replace the “greedy decode” in CoT prompting with sampling from the language model’s decoder to generate a diverse set of reasoning paths Wang et al. (2022). We apply temperature sampling Renze and Guven (2024) and for each input, there are $K$ pairs of rationales and corresponding labels.
@@ 3.2.2 Build positive and negative rationale sets
Language models are not perfect reasoners and they might occasionally produce an incorrect reasoning path or make a mistake in one of the reasoning steps. In general, evidence shows that correct reasoning processes, even though they are diverse, tend to have greater agreement in their final answers than incorrect reasoning processes Wang et al. (2022). We thus choose the rationales with the most consistent label in all outputs of the LLM as positive rationale samples, and the remaining rationales are viewed as negative rationale samples. However, the sampling times mainly influence the proportion of positive and negative samples where the number of negative samples is always smaller), along with the time and storage costs.
To deal with this issue, we conduct a self-play mechanism that the student model generates its own rationales by sampling from its previous iterations during the training (details can be seen in Appendix A.1), and we regard these low-quality rationales as negative samples based on the hypothesise that the rationale quality of LLMs is higher than that of smaller models. As a result, for each input question $\mathbf{x}=[x_{1},x_{2},...,x_{n}]$, we collect a positive rationale set $S_{pos}=\{\mathbf{r}_{1}^{pos},\mathbf{r}_{2}^{pos},...,\mathbf{r}_{m}^{pos}\}$ and a negative rationale set $S_{neg}=\{\mathbf{r}_{1}^{neg},\mathbf{r}_{2}^{neg},...,\mathbf{r}_{k}^{neg}\}$.
@@ 3.3 Contrastive knowledge distillation
In this subsection, we represents our designed quality guided contrastive rationale distillation for better knowledge learning from the perspective of training process.
@@ 3.3.1 Train a discriminator to judge rationales
The quality of rationales for the same question still differs. Moreover, as the training epoch increases, the rationales generated by the above self-play may become gradually closer to the positive rationales, and then viewing them as negative samples is no longer reasonable. For this reason, there is a need to train a discriminator $\mathcal{D}$ that can effectively judge the positive and negative rationales and output a score that represents the quality of each rationale. The input of the discriminator $\mathcal{D}$ is the question and the rationale, and we take an encoder architecture to measure the score, i.e.,
@@
We pretrain the $\mathcal{D}$ with the positive and negative rationales from the LLM, and during the training, the discriminator $\mathcal{D}$ is updated at regular epoch intervals (details can be seen in Appendix A.2), as the loss function can be formulated as
@@
Figure 4: Illustration of the MinMax scheme for contrastive distillation.
3.3.2 Quality guided contrastive distillation
@@
where $\mathbf{x}_{i}$, $S_{pos}^{{\color[rgb]{0,0,0}i}}$, and $S_{neg}^{{\color[rgb]{0,0,0}i}}$ denote the $i$-th question, and the corresponding positive and negative rationale sets, respectively, $N$ is the number of questions, $\beta>0$ is a tunable hyper-parameter, the function $l(\cdot)$ denotes the cross entropy loss, and the function $f(\cdot)$ denotes the rationale generation for its given input. Here,
@@
which are motivated by the sequence-level distillation Kim and Rush (2016) and illustrated in Fig. 4. Moreover, we set a margin $\delta$ for the negative rationales to balance the loss, i.e., $l(f(\mathbf{x}),\mathbf{r}_{j}^{neg})=\max(\delta-l(f(\mathbf{x}),\mathbf{r}_{j}^{neg}),0)$ with respect to the $j$-th negative rationale for an input question $\mathbf{x}$. Let us rethink the effectiveness of negative rationales generated by the previous-iteration smaller model, which enforces the smaller model to break out of local optima and yield a golden rationale that is closer to the output of the LLM. However, when the smaller model comes to converging, the previous-iteration smaller model is likely to output the rationales that are similar to those of the LLM, and then regarding them as negative samples is inaccurate. To address this issue, we introduce the quality guided distillation to optimize the training process and redefine the loss formulas in (4) and (5) as, respectively,
@@
where $s^{pos,i}$ and $s^{neg,i}$ are the corresponding quality scores obtained by the discriminator $\mathcal{D}$. It is reasonable that the positive rationales of higher quality should have larger weights across the datasets, while for the negative rationales of higher quality, it is on the contradiction. In the latter subsection 5.1, we will further discuss different schemes for the many-to-one distillation.

### v2

@@ Abstract
The deployment of large language models (LLMs) faces considerable challenges concerning resource constraints and inference efficiency. Recent research has increasingly focused on smaller, task-specific models enhanced by distilling knowledge from LLMs. However, prior studies have often overlooked the diversity and quality of knowledge, especially the untapped potential of negative knowledge. Constructing effective negative knowledge remains severely understudied. In this paper, we introduce a novel framework called quality-guided contrastive rationale distillation aimed at enhancing reasoning capabilities through contrastive knowledge learning. For positive knowledge, we enrich its diversity through temperature sampling and employ self-consistency for further denoising and refinement. For negative knowledge, we propose an innovative self-adversarial approach that generates low-quality rationales by sampling previous iterations of smaller language models, embracing the idea that one can learn from one’s own weaknesses. A contrastive loss is developed to distill both positive and negative knowledge into smaller language models, where an online-updating discriminator is integrated to assess qualities of rationales and assign them appropriate weights, optimizing the training process. Through extensive experiments across multiple reasoning tasks, we demonstrate that our method consistently outperforms existing distillation techniques, yielding higher-quality rationales. Our codes will be released soon.
QCRD: Quality-guided Contrastive Rationale Distillation for Large Language Models
Wei Wang1,2,  Zhaowei Li2,3,  Qi Xu2,  Yiqing Cai2,  Hang Song2,  Qi Qi2,  Ran Zhou2,  Zhida Huang2,  Tao Wang2,  Li Xiao1 1University of Science and Technology of China, 2ByteDance Inc, 3Fudan University wangweiii@mail.ustc.edu.cn
@@ 1 Introduction
The reasoning capabilities of large language models (LLMs) have been observed to scale their model sizes, while necessitating substantial memory and computing resources Chowdhery et al. (2023); Wei et al. (2022a). As such, efficient model compression is crucial in the deployment of LLMs, especially on resource-limited devices or platforms. Knowledge distillation from an LLM (teacher) to a smaller, more manageable language model (student) has recently emerged as a powerful and promising technique for model compression Hinton et al. (2015); Phuong and Lampert (2019). However, it is still open how to best reduce the performance gap between the teacher and the student on complex reasoning tasks Zelikman et al. (2022).
@@
In this regard, it has more recently been shown that adding explanation-augmented prompts, especially, Chain-of-Thought (CoT) Wei et al. (2022b), can enable LLMs to generate reasonable explanations (also referred to as rationales) to justify the reasoning outcomes Li et al. (2022). Distilling these rationales into smaller language models has been demonstrated to effectively improve the overall performance Hsieh et al. (2023); Li et al. (2022). For example, distilling Step-by-Step (DSS) Hsieh et al. (2023) was proposed as an innovative CoT distillation approach, which employed rationales from an LLM to guide a smaller language model under a multi-task learning setting. It involved training the smaller language model simultaneously on both label prediction and rationale generation tasks, effectively leveraging their mutual benefits.
@@
The essence of such distilling rationales is to guide the model in learning additional knowledge related to the labels. Knowledge can be generally concluded into two classes: positive and negative. Previous works on rationale distillation, although effective, still suffer from certain drawbacks. On the one hand, positive knowledge for distillation may be limited and noisy. Methods Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022); Chen et al. (2024b) treated rationales generated by LLMs as golden answers and aimed to minimize the gap between these rationales and those generated by smaller language models. However, despite LLMs’ powerful zero-shot/few-shot abilities, they may occasionally produce incorrect reasoning steps, leading to erroneous rationales/answers. Such erroneous rationales may degrade the reasoning performance of the distilled smaller language models. On the other hand, generating negative rationales and incorporating them into CoT distillation remain understudied, while negative knowledge has early proved constructive and effective for models.
@@
To this end, we propose a general method, named Quality-guided Contrastive Rationale Distillation (QCRD), to guide the knowledge distillation to smaller language models from a contrastive learning perspective. The comparison between the previous methods and our proposed QCRD is illustrated in Fig. 1. Specifically, the previous methods focus on the alignment between the rationale of the student model and the corresponding one of the teacher model, while our proposed QCRD aligns the student’s distribution and contrastive knowledge distribution with various sampled rationales. The core design of QCRD is to generate a diverse set of contrastive rationales and efficiently distill them into student models. For the positive part, to ensure the quality and variety of positive rationales, we prompt the LLM and sample the output to generate multi-round rationales for each input question. We then apply the self-consistency to denoise the rationale set and split it into positive rationales and negative rationales. For the negative part, we employ a self-adversarial strategy inspired by Silver et al. (2018) during training to generate low-quality rationales from previous iterations of smaller language models with a high sampling temperature and treat them as negative rationales. Finally, for better knowledge learning, we present a contrastive loss to distill both positive and negative rationales into smaller language models. A discriminator is adopted to assess the qualities of the rationales and assign them appropriate weights to optimize the training process across the datasets.
To demonstrate the superiority of QCRD, we conduct comprehensive experiments with two smaller types of T5 models Raffel et al. (2020), i.e., T5-base (220M parameters) and T5-small (60M parameters), on four popular datasets, followed by detailed analysis and discussion. Our main contributions of this paper can be summarized below.
@@
Figure 2: Illustration of the proposed quality-guided contrastive rationale distillation for distilling contrastive knowledge from teacher models into the student model. Fig.a represents our multi-task framework, i.e., the main prediction label task and additional rationale task. Fig.b represents generation of contrastive rationales for distillation. Fig.c represents details about the quality-guided contrastive rationale loss, and CE denotes the cross-entropy.
@@ 2 Related Work
Knowledge distillation from LLMs. Knowledge distillation (KD) is a highly effective technique for transferring knowledge from larger teacher models to smaller student models that are more suitable for practical applications Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022); Chen et al. (2024b); Wang et al. (2023). The KD technique can be generally classified into two different categories: (1) Black-box KD: only the teacher’s predictions are accessible; (2) White-box KD: it provides access to the teacher’s parameters. Both of them have shown promising potential in fine-tuning smaller models on the prompt response pairs generated by LLMs Zhu et al. (2023). In this paper, we hypothesize that only the predictions (predict labels and rationales) generated by LLMs are accessible.
Multi-task learning with LLM generated rationales. Current LLMs have already exhibited their capabilities to generate high-quality reasoning steps, resulting in rationales of their predictions Kojima et al. (2022), and these rationales have been found to be valuable additional knowledge for fine-tuning smaller models Hsieh et al. (2023). A multi-task learning framework is commonly employed that enforces smaller models to output corresponding rationales, while maintaining their original functionality. However, previous studies only focused on aligning the output of the smaller model with that of the LLM with a single loss form Hsieh et al. (2023); Magister et al. (2022).
Self-consistency of LLMs. The self-consistency of LLMs refers to the capacity to maintain coherent and rational reasoning during input processing. Based on the intuition that complex reasoning tasks typically admit multiple reasoning paths that reach a correct answer, the self-consistency can improve the LLMs’ reasoning performance by integratedly sampling CoT outputs several times and choosing the most consistent predict answer Stanovich and West (1991); Wang et al. (2022).
Contrastive learning for LLMs. Contrastive learning has demonstrated its efficiency across diverse domains, e.g., computer vision, natural language processing Jaiswal et al. (2020); Le-Khac et al. (2020). Notably, the application of contrastive learning to LLMs has recently emerged, highlighting the effectiveness of incorporating negative knowledge implicitly in model’s inputs and showing promising outcomes Li et al. (2024); Chen et al. (2024a). However, to the best of our knowledge, the application of contrastive learning in CoT rationale distillation has not been explored thus far.
@@ 3 Methodology
In this paper, we first propose a general contrastive CoT distillation approach, called quality-guided contrastive rationale distillation (QCRD), for training smaller models by distilling contrastive knowledge from teacher models. As illustrated in Fig. 2, our approach consists of the following three parts. (1) Following the method developed in Hsieh et al. (2023), we apply a multi-task learning framework for the supervised training of the student model, i.e., the main prediction label task and additional rationale generation task; see Fig. 2a. (2) As displayed in Fig. 2b, we design a general approach to generate contrastive knowledge from LLMs and student model itself for rationale distillation. (3) As shown in Fig. 2c, for better knowledge learning from rationales, we design a quality-guided contrastive learning strategy, where a contrastive loss is applied with the guidance of an online-updated discriminator to distinguish between positive and negative rationales and assign them quality scores.
3.1 Multi-task learning framework for the student model
Previous works have already demonstrated the advantages of the multi-task learning framework Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022). Accordingly, as shown in Fig. 2a, we apply the label prediction task and the rationale generation task to the training of smaller language models. Specifically, we use different prefixes to enforce smaller language models to generate different types of output. Given an input question, for the label prediction task, the smaller language model outputs the prediction label with input prefix $<Predict>$, while for the rationale generation task, it outputs the corresponding explanation with input prefix $<Explain>$. These outputs are then aligned to the corresponding ground truth and rationales using autoregressive loss, respectively.
@@ 3.2 Generation of contrastive knowledge
We use CoT prompting Wei et al. (2022b) to elicit and extract rationales from LLMs. As illustrated in Fig. 3, the LLM is provided with few examples to follow the output format. Instead of only generating one output for each input, we replace the “greedy decode” in CoT prompting with sampling from the language model’s decoder to generate a diverse set of reasoning paths Wang et al. (2022). We apply temperature sampling Renze and Guven (2024) to the LLM $K$ times, where a temperature value $\tau$ can control the diversity of the generated output. Therefore, for each input, there are $K$ pairs of rationales and corresponding labels.
@@ 3.2.1 Build positive and negative rationale sets
Language models are not perfect reasoners in the sense that they might occasionally produce an incorrect reasoning path or make a mistake in one of the reasoning steps. In general, evidence shows that correct reasoning processes, even though they are diverse, tend to have greater agreement in their final answers than incorrect reasoning processes Wang et al. (2022). We thus choose the rationales with the most consistent label in all outputs of the LLM as positive rationales, and the remaining rationales are viewed as negative rationales. However, the sampling times mainly influence the proportion of positive and negative samples (the number of negative samples is always smaller, and detailed statistical descriptions can be found in Appendix A.1), along with the time and storage costs. Moreover, these negative rationales from LLMs are likely to be positive for smaller language models, which may limit their effectiveness.
To deal with this issue, we conduct a self-adversarial mechanism that the student model generates its own negative rationales by sampling from its previous iterations with a high temperature value during training (we illustrate its rationality in Section 5.3 and demonstrate its superiority in Appendix A.2), and we regard these low-quality rationales as negative ones based on the hypothesis that the rationale quality of LLMs is higher than that of smaller models. As a result, for each input question $\mathbf{x}=[x_{1},x_{2},...,x_{n}]$, we collect a positive rationale set $S_{pos}=\{\mathbf{r}_{1}^{pos},\mathbf{r}_{2}^{pos},...,\mathbf{r}_{m}^{pos}\}$ and a negative rationale set $S_{neg}=\{\mathbf{r}_{1}^{neg},\mathbf{r}_{2}^{neg},...,\mathbf{r}_{k}^{neg}\}$.
@@ 3.3 Contrastive knowledge distillation
In this subsection, we present our designed quality-guided contrastive rationale distillation for better knowledge learning.
@@ 3.3.1 Train a discriminator to judge rationales
The quality of rationales for the same question still differs. Moreover, as the training epoch increases, the rationales generated by the above self-play may become gradually closer to the positive rationales, and then viewing them as negative ones is no longer reasonable. Therefore, there is a need to train a discriminator $\mathcal{D}$ that can effectively judge the positive and negative rationales and output a score that represents the quality of each rationale. The input of the discriminator $\mathcal{D}$ is the question and the rationale, and we take an encoder architecture to measure the score, i.e.,
@@
We pretrain the $\mathcal{D}$ with the positive and negative rationales from the LLM, and during training, the discriminator $\mathcal{D}$ is updated at regular epoch intervals (details can be seen in Appendix A.3). The loss function can be formulated as
@@
3.3.2 Quality-guided contrastive distillation
@@
where $\mathbf{x}_{i}$ denotes the $i$-th question, $S_{pos}^{i}$ and $S_{neg}^{i}$ denote the corresponding positive and negative rationale sets, respectively, $N$ is the number of questions, and $\beta>0$ is a tunable hyper-parameter. The function $l(\cdot)$ denotes the cross-entropy loss and $f(\cdot)$ denotes the rationale generation for its given input. In (3),
@@
which are motivated by the sequence-level distillation Kim and Rush (2016). Moreover, we set a margin $\delta$ for the negative rationales to filter out cases that are too simplistic, i.e., $l(f(\mathbf{x}),\mathbf{r}_{j}^{neg})=\min(l(f(\mathbf{x}),\mathbf{r}_{j}^{neg})-\delta,0)$ with respect to the $j$-th negative rationale for an input question $\mathbf{x}$. Let us rethink the effectiveness of negative rationales generated by the previous-iteration smaller model, which enforces the smaller model to break out of local optima and yield a golden rationale that is closer to the output of the LLM. However, when the smaller model comes to converging, the previous-iteration smaller model is likely to output the rationales that are similar to those of the LLM, and then regarding them as negative samples is inaccurate. To address this issue, we introduce the quality-guided distillation to optimize the training process and redefine the loss formulas in (4) and (5) as, respectively,
@@
where $s^{pos,i}$ and $s^{neg,i}$ are the corresponding quality scores obtained by the discriminator $\mathcal{D}$. By (6) and (7), the positive rationales of higher quality should have larger weights across the datasets, while for the negative rationales of higher quality, it is on the contradiction. In the latter sec. 5.1, we will further discuss different schemes for the many-to-one distillation.